Write the successor of:
 (b) 100199
(c) 1099999
(d) 2345670
 
egbjmu89erweghi84J9U85HUJITEJMH8I5TJ

Grade:12th pass

1 Answers

Harshit Singh
2 years ago
Dear Student

The successors are


(b) 100199 + 1 = 100200

(c) 1099999 + 1 = 1100000

​(d) 2345670 + 1 = 2345671

Thanks